# Towards Generalisable Time Series Understanding Across Domains

### v1

@@ Abstract
In natural language processing and computer vision, self-supervised pre-training on large datasets unlocks foundational model capabilities across domains and tasks. However, this potential has not yet been realised in time series analysis, where existing methods disregard the heterogeneous nature of time series characteristics. Time series are prevalent in many domains, including medicine, engineering, natural sciences, and finance, but their characteristics vary significantly in terms of variate count, inter-variate relationships, temporal dynamics, and sampling frequency. This inherent heterogeneity across domains prevents effective pre-training on large time series corpora. To address this issue, we introduce OTiS, an open model for general time series analysis, that has been specifically designed to handle multi-domain heterogeneity. We propose a novel pre-training paradigm including a tokeniser with learnable domain-specific signatures, a dual masking strategy to capture temporal causality, and a normalised cross-correlation loss to model long-range dependencies. Our model is pre-trained on a large corpus of $640,187$ samples and $11$ billion time points spanning $8$ distinct domains, enabling it to analyse time series from any (unseen) domain. In comprehensive experiments across $15$ diverse applications - including classification, regression, and forecasting - OTiS showcases its ability to accurately capture domain-specific data characteristics and demonstrates its competitiveness against state-of-the-art baselines. Our code and pre-trained weights are publicly available at https://github.com/oetu/otis.
@@ 1 Introduction
In natural language processing (NLP) or computer vision (CV), generalisable language features, e.g. semantics and grammar (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023), or visual features, e.g. edges and shapes (Geirhos et al., 2019; Dosovitskiy et al., 2021; Oquab et al., 2024), are learned from large-scale data. Self-supervised pre-training paradigms are designed to account for the specific properties of language (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023) or imaging (Zhou et al., 2022; Cherti et al., 2023; Oquab et al., 2024), unlocking foundational model capabilities that apply to a wide range of domains and downstream tasks. This potential, however, remains largely unrealised in time series due to the lack of self-supervised pre-training paradigms that account for the heterogeneity of time series across domains.
Time series are widespread in everyday applications and play an important role in various domains, including medicine (Pirkis et al., 2021), engineering (Gasparin et al., 2022), natural sciences (Ravuri et al., 2021), and finance (Sezer et al., 2020). They differ substantially with respect to the number of variates, inter-variate relationships, temporal dynamics, and sampling frequency (Fawaz et al., 2018; Ismail Fawaz et al., 2019; Ye & Dai, 2021; Wickstrøm et al., 2022). For instance, standard $10$-$20$ system electroencephalography (EEG) recordings come with up to $256$ variates (Jurcak et al., 2007), while most audio recordings have only $1$ (mono) or $2$ (stereo) variates. Weather data shows high periodicity, whereas financial data is exposed to long-term trends. Both domains encompass low-frequency data recorded on an hourly (278278278\,278mHz), daily ($12\,\mu$Hz), or even monthly (386386386\,386nHz) basis, while audio data is sampled at high frequencies of 44.144.144.1\,44.1kHz or more. Overall, this heterogeneity across domains renders the extraction of generalisable time series features difficult (Fawaz et al., 2018; Gupta et al., 2020; Iwana & Uchida, 2021; Ye & Dai, 2021).
While most existing self-supervised pre-training methods for time series are limited to single domains (Wu et al., 2021; 2022a; Nie et al., 2023; Dong et al., 2024; Jiang et al., 2024), recent works propose simple techniques to incorporate time series from multiple domains (Yang et al., 2024; Das et al., 2024; Woo et al., 2024; Liu et al., 2024). These works for instance crop all time series into segments of unified size (Jiang et al., 2024), resample them to a uniform frequency (Yang et al., 2024), or analyse each variate of a multi-variate time series independently (Liu et al., 2024). While these naive techniques address differences in sampling frequency and variate count, they degrade the original time series and neglect the critical inter-variate relationships and temporal dynamics required for effective real-world analysis. Consequently, there is a clear need for pre-training strategies that adequately handle heterogeneity in time series to unlock foundational model capabilities.
In this work, we propose a novel multi-domain pre-training paradigm that addresses the full spectrum of time series heterogeneity across domains. Our approach facilitates the comprehensive extraction of generalisable features from diverse time series. Pre-trained on a large corpus of publicly available data, our open model for general time series analysis (OTiS) can be fine-tuned on limited data of any (unseen) domain to perform a variety of downstream tasks, as showcased in Figure 1.
@@
Figure 1: Overview of OTiS. Pre-trained on a large corpus of time series from diverse domains, OTiS enables general time series analysis. Its domain-specific tokeniser addresses time series heterogeneity across domains - including different numbers of variates, inter-variate relationships, temporal dynamics, and sampling frequencies - by learning unique domain signatures. After pre-training, the model can be fine-tuned on limited data from any domain, including previously unseen ones, to perform various tasks such as classification, regression, and forecasting.
Our key contributions can be summarised as follows:
@@
    We present OTiS, an open model for general time series analysis, with our entire pipeline and pre-trained weights publicly available at https://github.com/oetu/otis.
@@
    We propose a novel pre-training paradigm based on masked data modelling to address heterogeneity in multi-domain time series. Our approach includes a novel tokeniser with learnable signatures to capture domain-specific data characteristics, a dual masking strategy to learn temporal causality, and a normalised cross-correlation loss to model long-range dependencies.
@@
    We pre-train OTiS on a large corpus of publicly available time series from $8$ domains, spanning medicine, engineering, natural sciences, and finance. With $640,187$ samples and $11$ billion time points, this corpus represents diverse time series characteristics, enabling generalisable feature extraction.
  4. 4.
    We evaluate OTiS across $15$ downstream applications, including classification, regression, and forecasting. Our comprehensive analysis demonstrates that OTiS accurately captures domain-specific data characteristics and is competitive with both specialised and general state-of-the-art (SOTA) models, achieving new SOTA performance in $10$ tasks. Notably, none of the baselines is capable of performing all the tasks covered by OTiS.
@@ 2.1 Self-Supervised Learning for Time Series
Time series vary significantly across domains, with differences in the number of variates, inter-variate relationships, temporal dynamics, and sampling frequencies. Due to this inherent heterogeneity, most existing works focus on pre-training models within a single domain (Oreshkin et al., 2019; Tang et al., 2020; Wu et al., 2021; Zhou et al., 2021; Wu et al., 2022a; Woo et al., 2022; Yue et al., 2022; Zhang et al., 2022; Li et al., 2023; Nie et al., 2023; Zeng et al., 2023; Dong et al., 2024). To develop more generalisable time series models, recent methods have explored multi-domain pre-training by addressing certain aspects of the heterogeneity, such as differences in variate count and sampling frequency. For instance, Liu et al. (2024) treat each variate in multi-variate time series independently to standardise generative tasks like forecasting, while Goswami et al. (2024) extend uni-variate analysis to discriminative tasks like classification. Similarly, Jiang et al. (2024) and Yang et al. (2024) standardise time series by cropping them into segments of predefined size and resampling them to a uniform frequency, respectively, to enable general classification capabilities in medical domains.
While partially addressing time series heterogeneity, these methods limit model capabilities for general time series analysis. Standardisation techniques like cropping or resampling may distort inter-variate relationships, temporal dynamics, and long-range dependencies. Additionally, many of these approaches are tailored to specific applications, such as generative tasks (Das et al., 2024; Liu et al., 2024; Woo et al., 2024), or focused on particular domains like medicine (Jiang et al., 2024; Yang et al., 2024). Moreover, recent foundational models (Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) focus on uni-variate analysis, ignoring crucial inter-variate relationships essential for real-world applications, such as disease prediction (Schoffelen & Gross, 2009; Wu et al., 2022b). Our study aims to overcome these limitations by fully addressing heterogeneity of multi-domain time series, establishing a foundation for general time series analysis across domains and tasks.
@@ 2.2 Time Series Tokenisation
Transformers (Vaswani et al., 2017) have emerged as the preferred architecture for foundational models in NLP and CV due to their scalability (Kaplan et al., 2020; Gordon et al., 2021; Alabdulmohsin et al., 2022), enabling the training of models in the magnitude of $100$ billion parameters (Chowdhery et al., 2023; Touvron et al., 2023; Oquab et al., 2024; Ravi et al., 2024). To utilise a Transformer for time series analysis, a tokeniser is required to map the time series into a compact latent space. Current methods (Jin et al., 2023; Nie et al., 2023; Zhou et al., 2023; Das et al., 2024; Goswami et al., 2024; Jiang et al., 2024; Liu et al., 2024; Woo et al., 2024; Yang et al., 2024) follow established techniques from NLP and CV, dividing time series into patches of pre-defined size. These patches are then flattened into a $1$D sequence, with positional embeddings used to retain positional information. While uni-variate models (Nie et al., 2023; Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) consider only temporal positions, multi-variate approaches (Woo et al., 2024; Yang et al., 2024; Jiang et al., 2024) account for both temporal and variate positions. However, none of these methods address the unique characteristics of variates, mistakenly assuming that the relationships between variates are identical across domains. Our work seeks to adapt the tokenisation process to preserve the domain-specific relationships between variates.
@@ 3 Methods
Figure 2: Architecture of OTiS. During pre-training, batches of time series from diverse domains are processed using a domain-specific tokeniser. This tokeniser splits a time series into fixed-size patches, which are then embedded using a patch projector shared across all variates and domains. A temporal embedding and a domain-specific variate embedding are added to each patch embedding. A dual masking strategy is employed to mask the resulting input tokens. The reconstruction of the multi-domain input tokens is guided using an auxiliary normalised cross-correlation (NCC) loss.
In this work, we present a novel multi-domain pre-training paradigm that enables generalisable feature extraction from large, heterogeneous time series corpora. We introduce a domain-specific tokeniser with learnable signatures to address heterogeneity in multi-domain time series, as described in Section 3.1. We tailor masked data modelling (MDM) for multi-domain time series to pre-train our open model for general time series analysis (OTiS) on a large, heterogeneous corpus, as detailed in Section 3.2. In particular, we propose normalised cross-correlation as a loss term to capture global temporal dynamics in time series, as explained in Section 3.3. Moreover, we introduce a dual masking strategy to capture bidirectional relationships and temporal causality, essential for general time series analysis, as described in Section 3.4. After pre-training, we fine-tune OTiS on limited data to perform a variety of downstream tasks in any - including previously unseen - domain, as outlined in Section 3.5. A graphical visualisation of our method is provided in Figure 2.
@@ 3.1 Domain-Specific Tokeniser
Assume a time series sample ${\bm{X}}\in\mathbb{R}^{V_{S}\times T}$ from domain $S$, where $V_{S}$ denotes the number of variates specific to $S$ and $T$ denotes the number of time points. We randomly crop or zero-pad ${\bm{X}}$ to a fixed context length of $\overline{T}$ time points. We then split it into $T^{\prime}$ temporal patches of size $P$ along the time dimension, resulting in $V_{S}\cdot T^{\prime}$ patches ${\bm{x}}_{v,t}\in\mathbb{R}^{1\times P}$, where $v\in\{1,\dots,V_{S}\}$ and $t\in\{1,\dots,T^{\prime}\}$.
Next, we embed these patches using a shared patch projector across all variates and domains, resulting in patch embeddings $e^{\mathcal{P}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}\in\mathbb{R}^{1\times D}$, where $D$ denotes the model dimension. The patch projector consists of a $1$D convolutional layer followed by layer normalisation and GELU activation.
@@
The permutation-equivariant nature of Transformers (Vaswani et al., 2017) requires the use of positional embeddings to accurately capture the inherent relationships in the input data. Initially introduced for $1$D textual token sequences (Vaswani et al., 2017), positional embeddings simply introduce an ordering into the input sequence. Modern implementations further extend their capabilities to encode more complex geometric information, such as $2$D spatial (Dosovitskiy et al., 2021) or graph (Kreuzer et al., 2021) structures. For the analysis of any-variate time series, we distinguish between the temporal and variate structure. The temporal structure is equivalent to a sequential $1$D structure, such that we use standard $1$D sinusoidal embeddings $e^{\mathcal{T}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{T}}_{t}\in\mathbb{R}^{1\times D}$.
@@
The variate structure exhibits great heterogeneity across domains. In domains with uni-variate and two-variate data, such as mono and stereo audio, the structure is either trivial or only requires a basic distinction between variates. In other domains, however, the variate structure may represent more complex relationships, such as $3$D manifolds for electroencephalography (EEG) or electrocardiography (ECG) data, or be of non-spatial nature, such as for financial data. Hence, we introduce learnable domain-specific variate embeddings to adequately address the heterogeneity across domains. These embeddings, denoted as $e^{\mathcal{V}}_{\mathcal{S}}(\displaystyle{\bm{x}}_{v,t})=\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$ for each variate $v$ in domain $S$, are designed to model the unique properties of a domain. They capture the inter-variate relationships and temporal dynamics specific to domain $S$, forming what can be considered as the signature of the very domain.
@@
Finally, the patch, temporal, and domain-specific variate embeddings are summed to form the input token ${\bm{e}}_{v,t}=\displaystyle{\bm{e}}^{\mathcal{P}}_{v,t}+\displaystyle{\bm{e}}^{\mathcal{T}}_{t}+\displaystyle{\bm{e}}^{\mathcal{V}}_{\mathcal{S},v}\in\mathbb{R}^{1\times D}$. These input tokens collectively constitute the final input sequence ${\bm{E}}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. To support batches of any-variate time series from multiple domains, we pad the variate dimension to the maximum number of variates in a batch $\overline{V}=\max_{S}V_{S}$. For samples where $V_{S}<\overline{V}$ or $T<\overline{T}$, attention masking is used to ensure that padded variate or temporal tokens are ignored. The domain-specific tokeniser is trained end-to-end with the Transformer layers.
Definition of (Sub-)Domains.
@@
The domain-specific tokeniser is designed to integrate different datasets within a domain. Consider two EEG datasets, TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), which share $19$ identical variates but have different sampling frequencies of 500500500\,500Hz and 200200200\,200Hz, respectively. In this case, a single EEG-specific tokeniser ($V_{\text{EEG}}=19$) is sufficient to accommodate both sampling frequencies, i.e. ${\bm{E}}^{\mathcal{V}}_{\text{EEG-TDBrain}}=\displaystyle{\bm{E}}^{\mathcal{V}}_{\text{EEG-SEED}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},19}]^{\top}\in\mathbb{R}^{19\times D}$, as demonstrated in our experiments in Section 4. Note that while these positional embeddings are agnostic to variate ordering, we simplify processing by aligning the variate order across datasets within the same domain. Consider another EEG dataset, LEMON (Babayan et al., 2019), which includes $62$ electrodes. Of these, $15$ overlap with the electrodes in TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), while the remaining $47$ are unique to LEMON (Babayan et al., 2019). In this scenario, the EEG-specific tokeniser can be extended by the $47$ new variates ($V_{\text{EEG}}=66$), such that ${\bm{E}}^{\mathcal{V}}_{\text{EEG-LEMON}}=[\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},1},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},15},\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},20},\dots,\displaystyle{\bm{e}}^{\mathcal{V}}_{\text{EEG},66}]^{\top}\in\mathbb{R}^{62\times D}$. In this way, different datasets can be combined to approximate the underlying data distribution of a domain $S$, e.g. EEG, enabling the creation of large and diverse time series corpora.

### v2

@@ Abstract
Recent breakthroughs in natural language processing and computer vision, driven by efficient pre-training on large datasets, have enabled foundation models to excel on a wide range of tasks. However, this potential has not yet been fully realised in time series analysis, as existing methods fail to address the heterogeneity in large time series corpora. Prevalent in domains ranging from medicine to finance, time series vary substantially in characteristics such as variate count, inter-variate relationships, temporal patterns, and sampling frequency. To address this, we introduce a novel pre-training paradigm specifically designed to handle time series heterogeneity. We propose a tokeniser with learnable domain signatures, a dual masking strategy, and a normalised cross-correlation loss, enabling our open model for general time series analysis (OTiS) to efficiently learn from large time series corpora. Extensive benchmarking on diverse tasks, such as classification, regression, and forecasting, demonstrates that OTiS outperforms state-of-the-art baselines. Our code and pre-trained weights are available at https://github.com/oetu/otis.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Open model for general time series analysis (OTiS). Its tokeniser accounts for varying time series characteristics across domains, such as distinct variate counts, inter-variate relationships, temporal patterns, and sampling frequencies. It can be fine-tuned on limited data from any domain, including previously unseen ones, to perform classification, regression, and forecasting tasks.
In natural language processing (NLP) or computer vision (CV), generalisable language features, e.g. semantics and grammar (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023), or visual features, e.g. edges and shapes (Geirhos et al., 2019; Dosovitskiy et al., 2021; Oquab et al., 2024), are learned from large-scale data. Self-supervised pre-training paradigms are designed to account for the specific properties of language (Radford et al., 2018; Touvron et al., 2023; Chowdhery et al., 2023) or imaging (Zhou et al., 2022; Cherti et al., 2023; Oquab et al., 2024), unlocking foundational model capabilities that apply to a wide range of tasks. This potential, however, remains largely unrealised in time series due to the lack of pre-training paradigms that effectively address time series heterogeneity across domains.
Time series are widespread in everyday applications and play an important role in various domains, including medicine (Pirkis et al., 2021), engineering (Gasparin et al., 2022), natural sciences (Ravuri et al., 2021), and finance (Sezer et al., 2020). They differ substantially with respect to the number of variates, inter-variate relationships, temporal patterns, and sampling frequency (Fawaz et al., 2018; Ismail Fawaz et al., 2019; Ye & Dai, 2021; Wickstrøm et al., 2022). For instance, standard $10$-$20$ system electroencephalography (EEG) recordings come with up to $256$ variates (Jurcak et al., 2007), while most audio recordings have only $1$ (mono) or $2$ (stereo) variates. Weather data shows high periodicity, whereas financial data is exposed to long-term trends. Both domains encompass low-frequency data recorded on an hourly ($278\,\mu$Hz), daily ($12\,\mu$Hz), or even monthly (386386386\,386nHz) basis, while audio data is sampled at high frequencies of 44.144.144.1\,44.1kHz or more. Overall, this heterogeneity across domains renders the extraction of generalisable time series features difficult (Fawaz et al., 2018; Gupta et al., 2020; Iwana & Uchida, 2021; Ye & Dai, 2021).
While most existing self-supervised pre-training methods for time series are limited to single domains (Wu et al., 2021, 2022a; Nie et al., 2023; Dong et al., 2024; Jiang et al., 2024), recent works propose simple techniques to incorporate time series from multiple domains (Yang et al., 2024; Das et al., 2024; Woo et al., 2024; Liu et al., 2024). These works, for instance, crop all time series into segments of unified size (Jiang et al., 2024), resample them to a uniform frequency (Yang et al., 2024), or analyse each variate of a multi-variate time series independently (Liu et al., 2024). While these naive techniques address domain differences in sampling frequency and variate count, they degrade the original time series and fail to account for inter-variate relationships and temporal patterns, which vary heavily across domains.
Consequently, there is a clear need for pre-training strategies that effectively address data heterogeneity across domains. In this work, we propose a novel multi-domain pre-training paradigm that accounts for the varying characteristics across domains. This approach enables our open model for general time series analysis (OTiS) to be pre-trained on a large and diverse time series corpus, unlocking foundational capabilities. Our key contributions can be summarised as follows:
@@
    We propose a novel multi-domain pre-training paradigm that enables the learning of generalisable time series features. We introduce a novel tokeniser to account for the varying characteristics across domains, a dual masking strategy to capture bidirectional relationships and temporal causality, and a normalised cross-correlation loss to learn global temporal patterns.
@@
    We pre-train OTiS on a large corpus spanning $8$ domains from medicine, engineering, natural sciences, and finance. With $640,187$ samples and $11$ billion time points, it covers diverse time series characteristics, essential for developing foundational model capabilities.
@@
    Evaluated across $15$ applications, OTiS is competitive with specialised and general state-of-the-art models, setting new benchmarks in $10$ tasks. Notably, none of the baselines can handle the variety of tasks covered by OTiS, i.e. classification, regression, and forecasting.
@@ 2.1 Self-Supervised Learning for Time Series
Time series vary significantly across domains, with differences in the number of variates, inter-variate relationships, temporal patterns, and sampling frequencies. Due to this inherent heterogeneity, most existing works focus on pre-training models within a single domain (Oreshkin et al., 2019; Tang et al., 2020; Wu et al., 2021; Zhou et al., 2021; Wu et al., 2022a; Woo et al., 2022; Yue et al., 2022; Zhang et al., 2022; Li et al., 2023; Nie et al., 2023; Zeng et al., 2023; Dong et al., 2024). To develop more general time series models, recent works have explored multi-domain pre-training by partially addressing this heterogeneity, including the different numbers of variates and sampling frequencies. Liu et al. (2024) treat each variate in a multi-variate time series independently to standardise generative tasks like forecasting, while Goswami et al. (2024) extend uni-variate analysis to discriminative tasks like classification. Jiang et al. (2024) and Yang et al. (2024) standardise time series by cropping them into segments of predefined size and resampling them to a uniform frequency, respectively, to enable general classification capabilities in medical domains.
While partially addressing time series heterogeneity, these pre-training paradigms cannot unlock foundational model capabilities. Standardisation techniques like cropping or resampling may distort inter-variate relationships and temporal patterns. Additionally, many of these approaches are tailored to specific applications, such as generative tasks (Das et al., 2024; Liu et al., 2024; Woo et al., 2024), or focus on domains in a particular field, like medicine (Jiang et al., 2024; Yang et al., 2024). Moreover, recent foundation models (Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) focus on uni-variate analysis, ignoring inter-variate relationships that are essential for real-world applications, such as disease prediction (Schoffelen & Gross, 2009; Wu et al., 2022b). Our study aims to overcome these limitations by effectively addressing time series heterogeneity, establishing a robust foundation for general time series analysis.
@@ 2.2 Time Series Tokenisation
Transformers (Vaswani et al., 2017) have emerged as the preferred architecture for foundation models in NLP and CV due to their scalability (Kaplan et al., 2020; Gordon et al., 2021; Alabdulmohsin et al., 2022), enabling the training of large models (Chowdhery et al., 2023; Touvron et al., 2023; Oquab et al., 2024; Ravi et al., 2024). To utilise a Transformer for time series analysis, a tokeniser is required to map the time series into a compact latent space. Current methods (Jin et al., 2023; Nie et al., 2023; Zhou et al., 2023; Das et al., 2024; Goswami et al., 2024; Jiang et al., 2024; Liu et al., 2024; Woo et al., 2024; Yang et al., 2024) follow established techniques from NLP and CV, dividing time series into patches of pre-defined size. These patches are then flattened into a $1$D sequence, with positional embeddings used to retain positional information. While uni-variate models (Nie et al., 2023; Das et al., 2024; Goswami et al., 2024; Liu et al., 2024) consider only temporal positions, multi-variate approaches (Woo et al., 2024; Yang et al., 2024; Jiang et al., 2024) account for both temporal and variate positions. However, none of the current methods consider the varying characteristics across domains, e.g. mistakenly assuming that the relationships between variates are identical across domains. Our work seeks to adapt time series tokenisation to account for these domain-specific characteristics.
Figure 2: Pre-training of OTiS. A time series is split into fixed-size patches, which are then embedded using a universal patch projector. In addition to a temporal embedding, each patch embedding is modulated with a domain-specific variate embedding to account for the unique characteristics of a domain. The resulting input tokens are masked with a dual masking strategy and reconstructed to optimise OTiS. The reconstruction is guided by a combination of mean squared error (MSE) and normalised cross-correlation (NCC) loss terms.
@@ 3 Methods
In this work, we present a novel multi-domain pre-training paradigm that enables the extraction of generalisable time series features. We introduce a tokeniser that accounts for varying time series characteristics across domains (see Section 3.1). We tailor masked data modelling to multi-domain time series, enabling our open model for general time series analysis (OTiS) to be pre-trained on a large, heterogeneous corpus (see Section 3.2). In particular, we introduce normalised cross-correlation as a loss term to capture global temporal patterns in time series (see Section 3.3) and propose a dual masking strategy to model both bidirectional relationships and temporal causality in time series (see Section 3.4). After pre-training, OTiS can be fine-tuned on limited data from any domain, including previously unseen ones, to perform general time series tasks, such as classification, regression, and forecasting (see Section 3.5).
@@ 3.1 Domain-Specific Tokeniser
Assume a time series sample $\bm{X}\in\mathbb{R}^{V_{S}\times T}$ from domain $S$, where $V_{S}$ denotes the number of variates specific to $S$ and $T$ denotes the number of time points. We randomly crop or zero-pad $\bm{X}$ to a fixed context length of $\overline{T}$ time points. We then split it into $T^{\prime}$ temporal patches of size $P$ along the time dimension, resulting in $V_{S}\cdot T^{\prime}$ patches $x_{v,t}\in\mathbb{R}^{1\times P}$, where $v\in\{1,\dots,V_{S}\}$ and $t\in\{1,\dots,T^{\prime}\}$.
Next, we embed these patches using a universal patch projector shared across domains, producing patch embeddings $e^{\mathcal{P}}(x_{v,t})=e^{\mathcal{P}}_{v,t}\in\mathbb{R}^{1\times D}$, where $D$ denotes the model dimension. The patch projector consists of a $1$D convolutional layer followed by layer normalisation and GELU activation. With a unified patch size and a domain-agnostic design, it serves as a general feature extractor and is inherently frequency-agnostic. While the sampling frequency varies across domains, it consistently ensures that relevant information is captured within the observation period, i.e. time series. Hence, as long as a model’s context length is sufficiently long, it has access to the relevant information within a time series, regardless of the sampling frequency.
@@
The permutation-equivariant nature of Transformers (Vaswani et al., 2017) requires the use of positional embeddings to accurately capture the inherent relationships of the input. Initially introduced for $1$D textual sequences (Vaswani et al., 2017), positional embeddings simply introduce an ordering into the input sequence. Modern implementations extend their capabilities to encode more complex geometric information, such as $2$D spatial (Dosovitskiy et al., 2021) or graph (Kreuzer et al., 2021) structures. In this work, we differentiate between temporal and variate structures for the analysis of any-variate time series. The temporal structure is equivalent to a $1$D sequence, such that we use standard $1$D sinusoidal embeddings $e^{\mathcal{T}}(x_{v,t})=e^{\mathcal{T}}_{t}\in\mathbb{R}^{1\times D}$.
@@
The variate structure exhibits great heterogeneity across domains. In domains with uni-variate and two-variate data, such as mono and stereo audio, the structure is either trivial or only requires a basic distinction between variates. In other domains, however, the variate structure may represent more complex relationships, such as $3$D manifolds for electroencephalography (EEG) or electrocardiography (ECG), or be of non-spatial nature, e.g. for financial data. To account for these varying characteristics across domains, we introduce learnable domain-specific variate embeddings. These embeddings, denoted as $e^{\mathcal{V}}_{S}(x_{v,t})=e^{\mathcal{V}}_{S,v}\in\mathbb{R}^{1\times D}$ for each variate $v$ in domain $S$, are designed to capture the unique characteristics of a domain $S$, including the domain-specific inter-variate relationships and temporal patterns. Hence, they can be considered as a domain signature after training.
@@
Finally, the patch, temporal, and domain-specific variate embeddings are summed to form the input token $e_{v,t}=e^{\mathcal{P}}_{v,t}+e^{\mathcal{T}}_{t}+e^{\mathcal{V}}_{S,v}\in\mathbb{R}^{1\times D}$. These tokens collectively form the final input sequence $\bm{E}\in\mathbb{R}^{(V_{S}\cdot T^{\prime})\times D}$. As time series across domains vary in the number of variates, we pad the variate dimension to the maximum number of variates $\overline{V}=\max_{S}V_{S}$ within each batch. For samples where $V_{S}<\overline{V}$ or $T<\overline{T}$, attention masking is used to ensure that padded variate or temporal tokens are ignored. The domain-specific tokeniser is trained end-to-end with the Transformer layers.
Definition of Domains.
@@
The domain-specific tokeniser is designed to integrate different datasets within a domain. Consider two EEG datasets, TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), which share $19$ identical variates but have different sampling frequencies of 500500500\,500Hz and 200200200\,200Hz, respectively. In this case, a single EEG-specific tokeniser ($V_{\text{EEG}}=19$) is sufficient to accommodate both sampling frequencies, i.e. $\bm{E}^{\mathcal{V}}_{\text{EEG-TDBrain}}=\bm{E}^{\mathcal{V}}_{\text{EEG-SEED}}=[e^{\mathcal{V}}_{\text{EEG},1},\dots,e^{\mathcal{V}}_{\text{EEG},19}]^{\top}\in\mathbb{R}^{19\times D}$, as shown in our experiments in Section 4. Note that while these positional embeddings are agnostic to variate ordering, we simplify processing by aligning the variate order across datasets within a domain. Consider another EEG dataset, LEMON (Babayan et al., 2019), with $62$ variates. Of these, $15$ overlap with the variates in TDBrain (Van Dijk et al., 2022) and SEED (Zheng & Lu, 2015), while the remaining $47$ are unique to LEMON (Babayan et al., 2019). In this scenario, the EEG-specific tokeniser can be extended by the $47$ new variates ($V_{\text{EEG}}=66$), such that $\bm{E}^{\mathcal{V}}_{\text{EEG-LEMON}}=[e^{\mathcal{V}}_{\text{EEG},1},\dots,e^{\mathcal{V}}_{\text{EEG},15},e^{\mathcal{V}}_{\text{EEG},20},\dots,e^{\mathcal{V}}_{\text{EEG},66}]^{\top}\in\mathbb{R}^{62\times D}$. In this way, different datasets can be combined to approximate the underlying data distribution of a domain, e.g. EEG, enabling the creation of large and diverse time series corpora.